Factory Overhead Rates and Account Balance

Prostheses Industries operates two factories. The manufacturing operations of Factory 1 are machine intensive, while the manufacturing operations of Factory 2 are labor intensive. The company applies factory overhead to jobs on the basis of machine hours in Factory 1 and on the basis of direct labor hours in Factory 2. Estimated factory overhead costs, direct labor hours, and machine hours are as follows:

Factory 1 Factory 2
Estimated factory overhead cost for fiscal
year beginning August 1 $699,300 $952,000
Estimated direct labor hours for year 14,000
Estimated machine hours for year 16,650
Actual factory overhead costs for August $55,960 $82,250
Actual direct labor hours for August 1,260
Actual machine hours for August 1,300

a. Determine the factory overhead rate for Factory 1. Round your answer to the nearest cent.
$ per machine hour

b. Determine the factory overhead rate for Factory 2. Round your answer to the nearest cent.
$ per direct labor hour

c. Determine the factory overhead applied to production in each factory for January.

Factory 1 $
Factory 2 $

d. Determine the balances of the factory accounts for each factory as of January 31, and indicate whether the amounts represent overapplied or underapplied factory overhead.

Factory 1 $
Factory 2 $

e. Explain why Factory 1 uses machine hours to allocate factory overhead while Factory 2 uses direct labor hours.
Factory overhead should be allocated using a base that is related to (causes) the overhead costs incurred. Factory 1 has a   manufacturing operation, and Factory 2 has a   manufacturing operation. Thus, Factory 1 uses machine hours, and Factory 2 uses direct labor hours to allocate factory overhead.
Answer #1

Answer:-

a) Factory overhead rate for factory 1 = 699,300/16,650 = 42 per machine hour

b) Factory overhead rate for factory 2 = 952,000/14,000 = 68 per labor hour

c) Calculate overhead applied :

Factory 1 = 1300*42 = 54,600

Factory 2 = 1,260*68 = 85,680

d) Calculate over and under applied overhead

Factory 1 = 54,600-55,960 = 690 Under applied

Factory 2 = 85,680-82,250 = 3,430 Over applied

if you have any query please ask me in comment box i am here to helps you don't give direct Thumbs down.if you satisfied my work give Thumbs UP

*******THANK YOU*******